# Positional encoding is not the same as context: A study on positional encoding for Sequential recommendation

### v1

@@ 4.3 Metrics
The second metric is Normalized Discounted Cumulative Gain (NDCG). NDCG considers the relevance of items and their positions in the recommended list.
$NDCG=\frac{DCG}{IDCG}$ (11)
@@
Where $DCG$ (Discounted Cumulative Gain) is calculated as the sum of the relevance scores of items at each position in the ranked list, discounted by their position, and $IDCG$ (Ideal Discounted Cumulative Gain) represents the maximum possible $DCG$ for a perfect ranking.
$DCG=\sum_{i=1}^{n}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$ (12)
$IDCG=\sum_{i=1}^{n}\frac{2^{max(rel_{i},0)}-1}{\log_{2}(i+1)}$ (13)
Here, $rel_{i}$ is the item’s relevance at position $i$ in the ranking.
5 Results
We present all the results for all the datasets and encodings in Appendix A.1. Due to space limitations, sub-tables were extracted from them.
Together with position encoding, we wanted to analyse other factors that could affect stability. In CARCA Rashed et al. (2022), they found $0.0001$ as the best upper bound ($nmax$ in our Tables) for the encoding vectors in all the datasets except for Games. We try different maximum values for all the encoding vectors ($0.1$ and $0.0001$) to see if their choice depends on the encoding. Except for some exceptions, the upper bound $0.0001$ produces better results. We also tried the activation function $silu$ instead of the $leakyrelu$. The results are slightly better regarding stability but without a strong correlation.
5.1 Stability
A common practice among the papers is to present an average of $5$. Our first experiments showed a significant difference between each one of the runs. We decided to analyse the deviation for the different trainings. The deviation and mean can be found in all the training tables. We also decided to compute the $95\%$ significance confidence interval and present it with the confidence interval in the complete tables results in Appendix A.1.
@@ 5.1 Stability
The stability varies from one dataset to another, being the Beauty dataset the most stable with an average Hit deviation of $0.9$ and NDCG deviation of $0.87$. Meanwhile, the maximum deviations are obtained by fashion. We can see that those datasets with a lower sparsity tend to be more stable, as do the Beauty and Games datasets, while those with a higher sparsity are less stable, as are the Men and Fashion datasets.
@@
In Table 6 and Table 12 in the Appendix, we see that some encodings produce more stable results than others. More precisely, the concatenated version for Abs , Learnt and Rotatory + Con. generally present more stable results. Furthermore, for the more stable case of $0.0001$ upper bound, in Table 6, Abs and Learnt are still significantly stable. At the same time, our Rotatory encoding produces similar results to the concatenated version. This could be because the average deviation is similar to no encoding, None . The most notable case is RMHA-4 , which, combined with the upper bound $0.0001$, gives the most stable results, around $0.51$, compared with the no-encoding case, $2.36$. This pattern is preserved to the general upper bound, suggesting that this encoding provides excellent stability. As we will see below, this encoding is worth trying in cases of high sparsity.
@@
We can appreciate that Learnt and the sinusoidal Abs tend to have a similar deviation, higher than not using any encoding. To know if the stability is provided by the fact that RMHA-4 introduce positional information in each transformer or being a RPE head, we trained Rope . The results show that it is a mid between RMHA-4 on the one hand and Learnt and Abs on the other hand. It does not add or reduce stability for None . We wanted to know if being Rope in a middle point is due to being in all the transformer blocks while not being a APE encoding. To this goal, we added RopeOne , an ablation which applies Rope but only in the first block. The results are not far away from each other. We can see in Table 12 that RopeOne is even slightly better, while in Table 6, they are almost the same, in this case, slightly worse. Then, we decided to create a new encoding, Rotatory , which uses the rotatory properties but is not in the attention heads but a vector encoding. The results show the same stability as Rope and its ablation.
The stability of the RMHA-4 encoding can also be appreciated in the losses. In Appendix A.5.
Dataset Hit Dev. NDCG Dev. CI-length Density/Sparsity-1
Men 3.9143 8.2742 5.6686 $56.727\times 10^{-}$ 5
Fashion 4.87 9.7814 9.2257 $54.765\times 10^{-}$ 5
Games 1.124 0.952 1.616 $43.904\times 10^{-}$ 4
Beauty 0.935 0.87 1.22 $41.320\times 10^{-}$ 4
Table 5: Deviations among the different datasets without Encoding. This pattern is generally preserved along the different encodings.
encoding dev HIT dev NDCG runs CI-length
Abs 5.43 8.25 9.38 6.71
Abs + Con 3.52 5.56 7.12 5.35
Learnt 4.79 7.64 7.88 7.05
Learnt + Con 3.17 6.36 6.62 4.42
None 2.35 5.26 8.00 3.29
RMHA-4 0.51 1.08 6.00 0.88
Rope 2.38 5.45 5.00 4.04
Rotatory 2.46 5.01 8.38 3.48
Rotatory + Con. 2.48 5.18 6.75 3.49
RopeOne 2.82 6.54 4.29 5.46
Rope-Longer 3.33 8.15 5.25 5.99
Rotatory-Longer 1.05 1.04 4.67 2.08
Rotatory + Con.-Longer 0.99 1.24 6.00 1.59
RMHA-4-Longer 0.71 4.00 6.00 1.14
Table 6: Deviations among the different encodings. Relative encoding in all the heads, RMHA-4 , presents more stability.
5.2 Best Results
@@
It depends on the spacing and the definition that one manages to get a good result. If we look at the table 7, we see that this dataset has excellent stability, but in cases like Learnt with $0.1$, with leakyrelu, the model gets over $74$ in HIT, but the $95\%$ confidence interval is ($60.74$, $88.24$). In other words, applying this approach cements the results of what was studied here Picard (2021); in some cases, it is better to get a good set of seeds than a suitable method.
@@
If we consider a maximum of $12$ as deviation, we get Table 3. In CARCA, they did an ablation study on the Men Dataset, showing that using Abs moves towards worse results. They produced similar results for Men, but both had a deviation of around $6$, so the results from the ablation were more a matter of initialisation. In this case, Men and Beauty get their best results with Learnt -like and Abs -like encodings, respectively. But they are unstable. In contrast, Fashion and Games get their top values with RMHA-4 and Rotatory , respectively, which are stable.
@@
If we reduce the maximum deviation to $3$, we get some results that we consider the right results. As we can see in Table 2, the best encodings are RMHA-4 and Rotatory (or its concatenated version), and they correlate with the sparsity. Those with a high sparsity benefit more from the extra stability provided by RMHA-4 , while those with a lower sparsity, Beauty and Games, get their best through a purely rotatory approach Rotatory . We can see that even considering only those which are stable, the results obtained in all the datasets are new SOTA.
Encoding "Ablation":
We created some ablations from existing encoding techniques to analyse where the stability came from. So far, our experiments used almost the same hyperparameters as the original CARCA. Looking again at the losses, we found that some of the top results did not converge, i.e., the loss and the metrics were improving. Thus, we decided to extend some trainings with more epochs to see if this produced better results. We added $400$ epochs to the default values (11). The improvements were not significant in this regard.
6 Conclusion
Our main conclusion is that encodings should not be considered minor in this model family. On the contrary, they are a powerful tool, and depending on the data set, a couple of possible better choices can be pre-selected. We have shown that the higher the sparsity, the better the results, as they stabilise the training until the best results are achieved. On the other hand, when the set density is sufficient, our rotatory encodings are a good choice or a learnable encoding. Also, at publication time, we have achieved SOTA results using only the encodings. And these results are, in most cases, more reliable than previous SOTA results.
7 Future work
During the encoding analysis, we found some correlation between the bound used for the encoding and the stability. Previous works have experimented with very low bounds, usually below $0.001$. We would like to go deep in this direction and analyse the behaviour of the metric used for encoding and its bounds.
8 Limitations
Despite the details of this study, it was carried out using a transformer architecture. Today, the latest models in the field are based on this architecture or GCNs. Using other architectures may affect this study and the conclusions on using encodings. The fact that the choice depends on the dataset can give a first idea of the possible encodings in case of using different architectures.
@@ Appendix A Appendix
A.1 Experiments
In this section, we present all our results for the four datasets, thereby completing the missing experiments from the sub-tables in the main corpus. In tables 7, 8, and 9, we can find Rope-Longer , which is not presented in Table 1. These are just a few cases in which we extend the number of epochs since we found that for Rope , it takes longer to reach its maximum.
Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
@@ A.1 Experiments
leaky Rotatory + Con. 0.0001 61.16 0.77 42.83 0.48 3 (60.29, 62.03) 1.74
silu Rotatory + Con. 0.0001 61.68 1.04 42.31 0.97 3 (60.5, 62.86) 2.36
leaky Rotatory + Con. 0.1000 50.51 1.59 32.25 1.76 3 (48.71, 52.31) 3.60
silu Rotatory + Con. 0.1000 50.88 1.00 32.70 0.87 3 (49.75, 52.01) 2.26
@@
Table 7: Beauty results with different positional encodings. Original results from Rashed et al. (2022) at the end.
Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
@@
leaky Rotatory + Con. 0.0001 69.13 2.80 55.33 7.69 9 (67.3, 70.96) 3.66
silu Rotatory + Con. 0.0001 71.16 1.38 48.61 8.62 7 (70.14, 72.18) 2.04
leaky Rotatory + Con. 0.1000 68.61 2.76 57.08 3.72 7 (66.57, 70.65) 4.08
silu Rotatory + Con. 0.1000 65.23 7.76 52.99 9.31 11 (60.64, 69.82) 9.18
@@
Table 8: Fashion Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
@@
leaky Rotatory + Con. 0.0001 70.64 5.29 59.30 6.98 10 (67.36, 73.92) 6.56
silu Rotatory + Con. 0.0001 67.18 5.82 50.27 13.21 10 (63.57, 70.79) 7.22
leaky Rotatory + Con. 0.1000 70.92 3.92 59.84 4.88 6 (67.78, 74.06) 6.28
silu Rotatory + Con. 0.1000 63.88 7.42 50.69 9.75 10 (59.28, 68.48) 9.20
@@
Table 9: Men Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
@@
leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79 6 (80.18, 81.06) 0.88
silu Rotatory + Con. NaN 79.61 2.17 54.80 2.71 6 (77.87, 81.35) 3.48
@@
leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93 6 (79.0, 81.42) 2.42
silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55 6 (79.91, 80.67) 0.76

### v2

@@ 4.3. Metrics
NDCG: Normalised Discounted Cumulative Gain evaluates the ranking quality by considering both the relevance of items and their positions in the recommended list. It is defined as:
(14) $NDCG=\frac{DCG}{IDCG}$
@@
where $DCG$ (Discounted Cumulative Gain) is calculated as the sum of the relevance scores of items at each position in the ranked list, discounted by their position:
(15) $DCG=\sum_{i=1}^{n}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$
and $IDCG$ (Ideal Discounted Cumulative Gain) represents the maximum possible $DCG$ for a perfectly ranked list:
(16) $IDCG=\sum_{i=1}^{n}\frac{2^{max(rel_{i},0)}-1}{\log_{2}(i+1)}$
Here, $rel_{i}$ denotes the relevance score of the item at position $i$ in the ranking. We will denote its mean as NDCG Mean, its deviation as NDCG Dev  and the average of the deviations as Avg Dev NDCG.
5. Results
We present comprehensive results for all datasets and encodings in Appendix A.1. Due to space limitations, only selected sub-tables are included in this section. Furthermore, Appendix A.8 presents an additional experiment using SASRec++, confirming the consistency of our results. Details about the resources used for these experiments are provided in Appendix A.3.
In addition to positional encoding, we analyze factors, such as maximum values and activation functions, that may influence model stability. In CARCA (Rashed et al., 2022), an upper bound of $0.0001$ ($nmax$ in our Tables) for the encoding vectors was identified as optimal across all datasets except for Games. To investigate whether the optimal upper bound varies with different encodings, we experimented with maximum values of $0.1$ and $0.0001$. Generally, an upper bound of $0.0001$ yielded better results, with a few exceptions. Furthermore, we compared activation functions by replacing $leakyrelu$ with $silu$. The $silu$ activation function resulted in slight improvements in stability; however, the correlation was not strong.
5.1. Stability
A common practice in the literature is to present the average results of $5$ runs with different random seeds. However, our preliminar experiments revealed significant variability between individual runs. To better understand this inconsistency, we analyzed the standard deviation across different trainings. Detailed deviation and mean values are provided in all the training tables. Additionally, we computed the $95\%$ confidence intervals for these metrics and included them in the complete results tables in Appendix A.1.
@@ 5.1. Stability
The stability of results varies across different datasets. Beauty is the most stable, showing an average Hit deviation of $0.9$ and an average NDCG deviation of $0.87$, while Fashion has the highest deviations. Datasets with lower sparsity, such as Beauty and Games, generally exhibit greater stability. In contrast, datasets with higher sparsity, like Men and Fashion, show reduced stability. Additionally, there is a correlation with data type: lower-sparsity datasets often use discrete embedding representations. As discussed in Section 5.3, this stability pattern persists even when we adjust sparsity levels. By reducing datasets to achieve comparable sparsity, we show that both encoding selection and stability challenges remain for each data type.
@@
Tables 4 and 19 (Appendix) reveal that some encodings yield more stable results than others. In particular, the concatenated versions of Abs , Learnt , and Rotatory demonstrate greater stability (Table 19). Under the most stable scenario with an upper bound of $0.0001$, as shown in Table 4, Abs + Con and Learnt + Con remain significantly more stable. However, Rotatory + Con shows no significant difference here, presenting deviations similar to None , likely due to the need to learn the angles.
The most notable encoding for stability is RMHA-4 , which, combined with the $0.0001$ upper bound, produces the most stable results, with an Avg Dev Hit  of $0.51$ compared to $2.36$ for the no-encoding case. This pattern persists even with higher upper bounds, suggesting that RMHA-4 offers excellent stability, which can also be observed in the loss graph, as detailed in Appendix A.6. As we will discuss in Section 5.3, this encoding is particularly worth considering in unstable scenarios.
@@
To assess whether the stability is due to RMHA-4  ’s positional information at each transformer block or its RPE  -based nature, we tested RoPE , an APE . The results rank RoPE between RMHA-4 and Learnt /Abs , showing no significant change in stability compared to None (Table 4). We hypothesized that this lack of change in stability could be due to RoPE ’s application across all transformer blocks. To test this, we introduced RopeOne  , which applies RoPE only in the first block. Results in Table 19 show that RopeOne slightly outperforms RoPE , while Table 4 shows similar performance with a slight lower stability. This suggests that positional information at each transformer block does not significantly impact stability. Finally, we introduced Rotatory , which incorporates rotatory properties in a vector encoding. The results showed similar stability to RoPE and its ablation RopeOne  . This led us to the conclusion that the rotatory nature of some encodings and the relative encoding nature of RMHA-4 contribute to the stability of the experiments.
5.2. Best Results
@@
The optimal encoding depends on the maximum deviation considered. For instance, examining the Beauty dataset in Table 9, we observe lower deviations. However, in cases like Learnt with $nmax=0.1$ and leakyrelu, despite achieving $74$ in Hit Mean, the Hit Dev  increases to $17.19$. This highlights the trade-off between performance and stability, as some encodings may offer better performance metrics at the cost of higher instability. These findings support the conclusions of (Picard, 2021), which suggest that, in certain scenarios, selecting a robust set of random seeds can be more important than choosing the best encoding method.
@@
When setting the maximum Hit Dev  to $12$, we get Table 6. In the CARCA (Rashed et al., 2022), an ablation analysis of the Men dataset revealed that using Abs results in poorer performance than None . Similar results were observed for the Men in our experiments, where both cases showed a deviation of around $6$, suggesting that the observed effects were more due to initialization than encoding choice. With a maximum of $12$ on Hit Dev, Men and Beauty achieve their best performance with Learnt -like and Abs -like encodings, respectively, although both are unstable. In contrast, Fashion and Games reach their optimal results with RMHA-4 and Rotatory , respectively, both of which are stable.
Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
Submen leaky RMHA-4 0.1000 77.28 2.66 68.45 3.65
Submen leaky RMHA-4 0.0001 75.96 2.71 66.03 4.00
Subfashion leaky RMHA-4 0.0001 79.72 2.85 71.57 3.66
Subfashion leaky None 0.1000 32.94 1.89 16.77 3.00
Subgames leaky Rotatory + Con NaN 52.23 1.57 31.51 0.96
Subgames silu Rotatory + Con NaN 50.73 0.89 30.48 0.72
Subgames leaky RMHA-4 NaN 50.43 0.95 29.96 0.87
Subgames leaky None NaN 49.53 1.01 28.84 0.69
Submen3 leaky Rotatory + Con 0.1000 73.12 2.38 62.86 2.99
Submen3 silu Rotatory + Con 0.0001 72.77 2.96 62.56 4.31
Table 7. The table displays the top $4$ results for each ablation dataset on cases with a Hit Devbelow $3$. In the case of the Fashion dataset ablation we did not get any training below $3$ in the deviation.
@@
Reducing the maximum Hit Dev  to $3$ yields optimal results, as presented in Table 5. In this Table, the best encodings are RMHA-4 and Rotatory (or its concatenated version), which correlate strongly with sparsity, datatype, and stability (low-deviation). As discussed in Section 5.3, datasets with high-deviation without encoding benefit from the stability provided by RMHA-4 , while low-deviation datasets like Beauty and Video Games perform best with the purely rotatory approach of Rotatory . Even among stable datasets, these results surpass the previous SOTA benchmarks (CARCA) at the time. We also observe a significant improvement over the None case for high-deviation datasets, with gains exceeding $5\%$, while achieving greater stability.
Impact of Extended Training on Encoding Stability:
Ablation studies were conducted on the encoding techniques to investigate the source of stability. Our initial experiments used hyperparameters similar to those of the original CARCA model (Appendix A.4). Upon reviewing the loss metrics, we found that several top-performing models did not converge, as both loss and performance metrics continued to improve. To address this, we extended the training by an additional $400$ epochs beyond the default values specified in Table 18. Despite this extension, the improvements in model performance were minimal, indicating that the optimizer may have been oscillating around a local minimum.
5.3. Sparsity and Deviation Reduction
As seen in Section 5.2, two encodings, RMHA-4 and Rotatory + Con , consistently provide the best results (with maximum Hit Dev  of $3$). There appears to be a correlation between the deviation of datasets and the choice of optimal encoding. For high-deviation datasets like Men and Fashion, RMHA-4 yields better results, whereas for low-deviation datasets like Games and Beauty, Rotatory + Con performs best. Importantly, both Men and Fashion share two other key characteristics: dense vector representations and higher-sparsity. This makes it difficult to attribute the performance solely to the default dataset deviation. The complete tables for the subset results are available in Appendix A.2.
Sparsity Reduction:
To investigate this further, We conducted a sparsity reduction experiment by creating subdatasets with densities comparable to other datasets. Specifically, we reduced the sparsity of Fashion from $4.765\cdot 10^{-5}$ to $3.564\cdot 10^{-4}$ and Men from $6.727\cdot 10^{-5}$ to $2.860\cdot 10^{-4}$, resulting in Subfashion and Submen. We applied RMHA-4 , Rotatory + Con , and None to these datasets.
Table 7 shows that RMHA-4 consistently produced the best results across both subdatasets, confirming that its effectiveness is not solely due to sparsity. In Subfashion, Hit Devincreased to $15.105$, and NDCG Devto $17.935$, while Submen experienced a small reduction in deviations (Hit Dev7.2967.2967.2967.296, NDCG Dev9.0969.0969.0969.096). Although Men saw a slight reduction on deviation (Table 8), it wasn’t substantial.
Additionally, we created Subgames for comparison, reducing the Games dataset’s sparsity to see if it influences performance. Interestingly, Rotatory + Con remained the best encoding for this more stable dataset (Table 7), suggesting that even after sparsity reduction, dataset characteristics like deviation may drive encoding performance.
Deviation Reduction:
To rule out the correlation with the vector representation and isolate the effect of deviation, we further reduced the Men dataset to $10k$ users and $80k$ items ($10000$ users and $45129$ items after filtering). This reduced sparsity to $1.648\cdot 10^{-4}$ and helped stabilize the experiments, decreasing Hit Devto $5.198$ and NDCG Devto $6.826$. The reduced deviation was enough to make Rotatory + Con the best-performing encoding. Although the performance is close to that of the None case (Table 17), there was a shift in the optimal encoding.
These findings indicate that high-deviation datasets, even after sparsity reduction, benefit from the added stability of RMHA-4 . In contrast, low-deviation datasets like Subgames, perform better with Rotatory + Con  .
Dataset Users Items Interac. Hit Dev NDCG Dev CI-length Density
Subfashion 3,840 5,000 16,080 15.105 17.935 24.175 $3.564\cdot 10^{-4}$
Submen 8,332 10,000 28,603 7.2925 9.0925 10.515 $2.860\cdot 10^{-4}$
Subgames 5,000 4,601 33,353 1.115 0.775 1.79 $1.3341\cdot 10^{-3}$
Submen 2 10,000 45,129 74,391 5.1975 6.8225 8.31 $1.648\cdot 10^{-4}$
Table 8. Statistics of the ablations: Deviations for Hit@10 and NDCG without Encoding. Together with the average length of the $CI$ intervals, and the density.
6. Conclusion
In this study, we emphasize that encodings are not mere technical details but rather critical components that can significantly influence both the performance and stability of SRS. Our findings reveal that high-deviation datasets benefit from relative encodings, such as RMHA-4 , which not only stabilize training but also yield optimal results. These encodings play a vital role in mitigating performance fluctuations, allowing the model to consistently perform well across different runs. On the other hand, for low-deviation datasets, our proposed Rotatory encoding offers a robust alternative that further improves results without introducing unnecessary complexity.
Importantly, at the time of publication, we achieved state-of-the-art performance using only these encoding methods, surpassing previous benchmarks and demonstrating greater reliability. Our results indicate that encodings are a pivotal factor in ensuring the stability of transformer-based models, especially in recommendation systems. We encourage the research community to place greater emphasis on positional encodings, as they can drastically enhance model effectiveness, particularly in tasks involving sequential data.
In the Appendix A.7, there is a decision tree for the encoding election.
@@ Appendix A Appendix
A.1. Experiments
In this section, we present all our results for the four datasets, thereby completing the missing experiments from the sub-tables in the main corpus. In tables 9, 10, and 11, we can find RoPE-Longer , which is not presented in Table 1. These are just a few cases in which we extend the number of epochs since we found that for RoPE , it takes longer to reach its maximum.
Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
@@ A.1. Experiments
leaky Rotatory + Con 0.0001 61.16 0.77 42.83 0.48 3 (60.29, 62.03) 1.74
silu Rotatory + Con 0.0001 61.68 1.04 42.31 0.97 3 (60.5, 62.86) 2.36
leaky Rotatory + Con 0.1000 50.51 1.59 32.25 1.76 3 (48.71, 52.31) 3.60
silu Rotatory + Con 0.1000 50.88 1.00 32.70 0.87 3 (49.75, 52.01) 2.26
@@
Table 9. Beauty results with different positional encodings. Original results from (Rashed et al., 2022) at the end.
Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
@@
leaky Rotatory + Con 0.0001 69.13 2.80 55.33 7.69 9 (67.3, 70.96) 3.66
silu Rotatory + Con 0.0001 71.16 1.38 48.61 8.62 7 (70.14, 72.18) 2.04
leaky Rotatory + Con 0.1000 68.61 2.76 57.08 3.72 7 (66.57, 70.65) 4.08
silu Rotatory + Con 0.1000 65.23 7.76 52.99 9.31 11 (60.64, 69.82) 9.18
@@
Table 10. Fashion Dataset results with different positional encodings. Results from (Rashed et al., 2022) at the end.
Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
@@
leaky Rotatory + Con 0.0001 70.64 5.29 59.30 6.98 10 (67.36, 73.92) 6.56
silu Rotatory + Con 0.0001 67.18 5.82 50.27 13.21 10 (63.57, 70.79) 7.22
leaky Rotatory + Con 0.1000 70.92 3.92 59.84 4.88 6 (67.78, 74.06) 6.28
silu Rotatory + Con 0.1000 63.88 7.42 50.69 9.75 10 (59.28, 68.48) 9.20
@@
Table 11. Men Dataset results with different positional encodings. Results from (Rashed et al., 2022) at the end.
Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
@@
leaky Rotatory + Con NaN 80.62 0.55 56.07 0.79 6 (80.18, 81.06) 0.88
silu Rotatory + Con NaN 79.61 2.17 54.80 2.71 6 (77.87, 81.35) 3.48
@@
leaky Rotatory + Con-Longer NaN 80.21 1.51 55.06 1.93 6 (79.0, 81.42) 2.42
silu Rotatory + Con-Longer NaN 80.29 0.47 55.18 0.55 6 (79.91, 80.67) 0.76